OpenStudy (brunsonni):

David kicks a soccer ball off the ground and in the air, with an initial velocity of 35 feet per second. Using the formula H(t)=-16t^2+vt+s, what is the maximum height the soccer ball reaches? A. 17.9 feet B. 18.2 feet C. 18.7 feet D. 19.1 feet Please help, I have more than one question like this and it's a module test, so I can't get it wrong. I don't understand how to do this at all.

OpenStudy (anonymous):

v is supposed to be the initial speed of 35 and s is the initial height which is 0, so the equation is \[h(t)=-16t^2+35t\], but that won't work with your options.

OpenStudy (welshfella):

at height 0 -16t^2 - 35t = 0 - we can convert this to vertex form and find maximum height

OpenStudy (brunsonni):

@peachpi I still don't understand what I would do after getting that equation, because the one I had was the exact same but I had + s on the end. But what would I do next after getting h(t)=-16t^2+35t?

OpenStudy (welshfella):

* -16t^2 + 35t = 0

OpenStudy (brunsonni):

@welshfella would I do -16t^2=0 and -35t=0 now or is that completely wrong

OpenStudy (anonymous):

Hi @brunsonnie use this, this might help https://mathway.com/

OpenStudy (anonymous):

You could do t = -b/(2a) to find the time of the max height. t = -35/(2*-16) Then plug that in for the height

OpenStudy (welshfella):

take out -16 -16[t^2 + 2.1875] = -16[(x - 1.09375)^2 - 1.1964] maxm height = 16 * 1.1964 = 19.14

OpenStudy (welshfella):

* thats t not x!

OpenStudy (welshfella):

the 1.1964 comes fro 1.09375^2

OpenStudy (brunsonni):

Thanks

OpenStudy (welshfella):

yw